\begin{thebibliography}{28}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Ba{\~n}{\'o}n et~al.(2020)Ba{\~n}{\'o}n, Chen, Haddow, Heafield,
  Hoang, Espl{\`a}-Gomis, Forcada, Kamran, Kirefu, Koehn, Ortiz~Rojas,
  Pla~Sempere, Ram{\'\i}rez-S{\'a}nchez, Sarr{\'\i}as, Strelec, Thompson,
  Waites, Wiggins, and Zaragoza}]{banon-etal-2020-paracrawl}
Marta Ba{\~n}{\'o}n, Pinzhen Chen, Barry Haddow, Kenneth Heafield, Hieu Hoang,
  Miquel Espl{\`a}-Gomis, Mikel~L. Forcada, Amir Kamran, Faheem Kirefu, Philipp
  Koehn, Sergio Ortiz~Rojas, Leopoldo Pla~Sempere, Gema
  Ram{\'\i}rez-S{\'a}nchez, Elsa Sarr{\'\i}as, Marek Strelec, Brian Thompson,
  William Waites, Dion Wiggins, and Jaume Zaragoza. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.417} {{P}ara{C}rawl:
  Web-scale acquisition of parallel corpora}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 4555--4567, Online. Association for
  Computational Linguistics.

\bibitem[{Bao et~al.(2021)Bao, Zhang, Teng, Chen, and Luo}]{bao-etal-2021-g}
Guangsheng Bao, Yue Zhang, Zhiyang Teng, Boxing Chen, and Weihua Luo. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.267}
  {{G}-transformer for document-level machine translation}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 3442--3455,
  Online. Association for Computational Linguistics.

\bibitem[{Bawden et~al.(2018)Bawden, Sennrich, Birch, and
  Haddow}]{bawden-etal-2018-evaluating}
Rachel Bawden, Rico Sennrich, Alexandra Birch, and Barry Haddow. 2018.
\newblock \href {https://doi.org/10.18653/v1/N18-1118} {Evaluating discourse
  phenomena in neural machine translation}.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, pages 1304--1313, New Orleans,
  Louisiana. Association for Computational Linguistics.

\bibitem[{Donato et~al.(2021)Donato, Yu, and Dyer}]{donato-etal-2021-diverse}
Domenic Donato, Lei Yu, and Chris Dyer. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.104} {Diverse
  pretrained context encodings improve document translation}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 1299--1311,
  Online. Association for Computational Linguistics.

\bibitem[{Fernandes et~al.(2021)Fernandes, Yin, Neubig, and
  Martins}]{fernandes-etal-2021-measuring}
Patrick Fernandes, Kayo Yin, Graham Neubig, and Andr{\'e} F.~T. Martins. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.505} {Measuring and
  increasing context usage in context-aware machine translation}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 6467--6478,
  Online. Association for Computational Linguistics.

\bibitem[{Jauregi~Unanue et~al.(2020)Jauregi~Unanue, Esmaili, Haffari, and
  Piccardi}]{jauregi-unanue-etal-2020-leveraging}
Inigo Jauregi~Unanue, Nazanin Esmaili, Gholamreza Haffari, and Massimo
  Piccardi. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.coling-main.395} {Leveraging
  discourse rewards for document-level neural machine translation}.
\newblock In \emph{Proceedings of the 28th International Conference on
  Computational Linguistics}, pages 4467--4482, Barcelona, Spain (Online).
  International Committee on Computational Linguistics.

\bibitem[{Kang et~al.(2020)Kang, Zhao, Zhang, and
  Zong}]{kang-etal-2020-dynamic}
Xiaomian Kang, Yang Zhao, Jiajun Zhang, and Chengqing Zong. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.175} {Dynamic
  context selection for document-level neural machine translation via
  reinforcement learning}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 2242--2254, Online. Association
  for Computational Linguistics.

\bibitem[{Koehn(2004)}]{koehn-2004-statistical}
Philipp Koehn. 2004.
\newblock \href {https://aclanthology.org/W04-3250} {Statistical significance
  tests for machine translation evaluation}.
\newblock In \emph{Proceedings of the 2004 Conference on Empirical Methods in
  Natural Language Processing}, pages 388--395, Barcelona, Spain. Association
  for Computational Linguistics.

\bibitem[{Koehn et~al.(2007)Koehn, Hoang, Birch, Callison-Burch, Federico,
  Bertoldi, Cowan, Shen, Moran, Zens, Dyer, Bojar, Constantin, and
  Herbst}]{koehn-etal-2007-moses}
Philipp Koehn, Hieu Hoang, Alexandra Birch, Chris Callison-Burch, Marcello
  Federico, Nicola Bertoldi, Brooke Cowan, Wade Shen, Christine Moran, Richard
  Zens, Chris Dyer, Ond{\v{r}}ej Bojar, Alexandra Constantin, and Evan Herbst.
  2007.
\newblock \href {https://aclanthology.org/P07-2045} {{M}oses: Open source
  toolkit for statistical machine translation}.
\newblock In \emph{Proceedings of the 45th Annual Meeting of the Association
  for Computational Linguistics Companion Volume Proceedings of the Demo and
  Poster Sessions}, pages 177--180, Prague, Czech Republic. Association for
  Computational Linguistics.

\bibitem[{Kuang and Xiong(2018)}]{kuang-xiong-2018-fusing}
Shaohui Kuang and Deyi Xiong. 2018.
\newblock \href {https://aclanthology.org/C18-1051} {Fusing recency into neural
  machine translation with an inter-sentence gate model}.
\newblock In \emph{Proceedings of the 27th International Conference on
  Computational Linguistics}, pages 607--617, Santa Fe, New Mexico, USA.
  Association for Computational Linguistics.

\bibitem[{Kuang et~al.(2018)Kuang, Xiong, Luo, and
  Zhou}]{kuang-etal-2018-modeling}
Shaohui Kuang, Deyi Xiong, Weihua Luo, and Guodong Zhou. 2018.
\newblock \href {https://aclanthology.org/C18-1050} {Modeling coherence for
  neural machine translation with dynamic and topic caches}.
\newblock In \emph{Proceedings of the 27th International Conference on
  Computational Linguistics}, pages 596--606, Santa Fe, New Mexico, USA.
  Association for Computational Linguistics.

\bibitem[{Ma et~al.(2020)Ma, Zhang, and Zhou}]{ma-etal-2020-simple}
Shuming Ma, Dongdong Zhang, and Ming Zhou. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.321} {A simple and
  effective unified encoder for document-level machine translation}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 3505--3511, Online. Association for
  Computational Linguistics.

\bibitem[{Maruf and Haffari(2018)}]{maruf-haffari-2018-document}
Sameen Maruf and Gholamreza Haffari. 2018.
\newblock \href {https://doi.org/10.18653/v1/P18-1118} {Document context neural
  machine translation with memory networks}.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1275--1284,
  Melbourne, Australia. Association for Computational Linguistics.

\bibitem[{Maruf et~al.(2019)Maruf, Martins, and
  Haffari}]{maruf-etal-2019-selective}
Sameen Maruf, Andr{\'e} F.~T. Martins, and Gholamreza Haffari. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1313} {Selective attention for
  context-aware neural machine translation}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 3092--3102,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{M{\"u}ller et~al.(2018)M{\"u}ller, Rios, Voita, and
  Sennrich}]{muller-etal-2018-large}
Mathias M{\"u}ller, Annette Rios, Elena Voita, and Rico Sennrich. 2018.
\newblock \href {https://doi.org/10.18653/v1/W18-6307} {A large-scale test set
  for the evaluation of context-aware pronoun translation in neural machine
  translation}.
\newblock In \emph{Proceedings of the Third Conference on Machine Translation:
  Research Papers}, pages 61--72, Brussels, Belgium. Association for
  Computational Linguistics.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and
  Zhu}]{papineni2002bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock Bleu: a method for automatic evaluation of machine translation.
\newblock In \emph{Proceedings of the 40th annual meeting of the Association
  for Computational Linguistics}, pages 311--318.

\bibitem[{Rei et~al.(2020)Rei, Stewart, Farinha, and
  Lavie}]{rei-etal-2020-comet}
Ricardo Rei, Craig Stewart, Ana~C Farinha, and Alon Lavie. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.213} {{COMET}: A
  neural framework for {MT} evaluation}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 2685--2702, Online. Association
  for Computational Linguistics.

\bibitem[{Saunders et~al.(2020)Saunders, Stahlberg, and
  Byrne}]{saunders-etal-2020-using}
Danielle Saunders, Felix Stahlberg, and Bill Byrne. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.693} {Using context
  in neural machine translation training objectives}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 7764--7770, Online. Association for
  Computational Linguistics.

\bibitem[{Sennrich et~al.(2016)Sennrich, Haddow, and
  Birch}]{sennrich-etal-2016-neural}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016.
\newblock \href {https://doi.org/10.18653/v1/P16-1162} {Neural machine
  translation of rare words with subword units}.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1715--1725,
  Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Tu et~al.(2018)Tu, Liu, Shi, and Zhang}]{tu-etal-2018-learning}
Zhaopeng Tu, Yang Liu, Shuming Shi, and Tong Zhang. 2018.
\newblock \href {https://doi.org/10.1162/tacl_a_00029} {Learning to remember
  translation history with a continuous cache}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  6:407--420.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock Attention is all you need.
\newblock \emph{Advances in neural information processing systems}, 30.

\bibitem[{Voita et~al.(2019)Voita, Sennrich, and Titov}]{voita-etal-2019-good}
Elena Voita, Rico Sennrich, and Ivan Titov. 2019.
\newblock \href {https://doi.org/10.18653/v1/P19-1116} {When a good translation
  is wrong in context: Context-aware machine translation improves on deixis,
  ellipsis, and lexical cohesion}.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 1198--1212, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Voita et~al.(2018)Voita, Serdyukov, Sennrich, and
  Titov}]{voita-etal-2018-context}
Elena Voita, Pavel Serdyukov, Rico Sennrich, and Ivan Titov. 2018.
\newblock \href {https://doi.org/10.18653/v1/P18-1117} {Context-aware neural
  machine translation learns anaphora resolution}.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1264--1274,
  Melbourne, Australia. Association for Computational Linguistics.

\bibitem[{Xu et~al.(2021)Xu, Li, Wong, Liu, and
  Chao}]{xu-etal-2021-document-graph}
Mingzhou Xu, Liangyou Li, Derek~F. Wong, Qun Liu, and Lidia~S. Chao. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.emnlp-main.663} {Document
  graph for neural machine translation}.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, pages 8435--8448, Online and Punta Cana,
  Dominican Republic. Association for Computational Linguistics.

\bibitem[{Yang et~al.(2019)Yang, Zhang, Meng, Gu, Feng, and
  Zhou}]{yang-etal-2019-enhancing}
Zhengxin Yang, Jinchao Zhang, Fandong Meng, Shuhao Gu, Yang Feng, and Jie Zhou.
  2019.
\newblock \href {https://doi.org/10.18653/v1/D19-1164} {Enhancing context
  modeling with a query-guided capsule network for document-level translation}.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pages 1527--1537, Hong Kong,
  China. Association for Computational Linguistics.

\bibitem[{Yu et~al.(2020)Yu, Sartran, Stokowiec, Ling, Kong, Blunsom, and
  Dyer}]{yu-etal-2020-better}
Lei Yu, Laurent Sartran, Wojciech Stokowiec, Wang Ling, Lingpeng Kong, Phil
  Blunsom, and Chris Dyer. 2020.
\newblock \href {https://doi.org/10.1162/tacl_a_00319} {Better document-level
  machine translation with {B}ayes{'} rule}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  8:346--360.

\bibitem[{Zhang et~al.(2018)Zhang, Luan, Sun, Zhai, Xu, Zhang, and
  Liu}]{zhang-etal-2018-improving}
Jiacheng Zhang, Huanbo Luan, Maosong Sun, Feifei Zhai, Jingfang Xu, Min Zhang,
  and Yang Liu. 2018.
\newblock \href {https://doi.org/10.18653/v1/D18-1049} {Improving the
  transformer translation model with document-level context}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 533--542, Brussels, Belgium. Association
  for Computational Linguistics.

\bibitem[{Zheng et~al.(2020)Zheng, Yue, Huang, Chen, and Birch}]{ijcai2020p551}
Zaixiang Zheng, Xiang Yue, Shujian Huang, Jiajun Chen, and Alexandra Birch.
  2020.
\newblock \href {https://doi.org/10.24963/ijcai.2020/551} {Towards making the
  most of context in neural machine translation}.
\newblock In \emph{Proceedings of the Twenty-Ninth International Joint
  Conference on Artificial Intelligence, {IJCAI-20}}, pages 3983--3989.
  International Joint Conferences on Artificial Intelligence Organization.
\newblock Main track.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{EACL2023}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\title{Exploring    Paracrawl for Document-level Neural Machine Translation}

\author{Yusser Al Ghussin$^{1,2}$, Jingyi Zhang$^3$ and Josef van Genabith$^{1,2}$ \\
  $^1$German Research Center for Artificial Intelligence (DFKI),\\
Saarland Informatics Campus, Saarbrucken, Germany  \\
$^2$Department of Language Science and Technology, Saarland University, Germany\\
  $^3$Hasso-Plattner-Institut (HPI),  Potsdam, Germany\\
  \texttt{yusser.al\_ghussin/Josef.Van\_Genabith@dfki.de,Jingyi.Zhang@hpi.de} \\}

\begin{document}
\maketitle
\begin{abstract}

 Document-level neural machine translation (NMT) has outperformed sentence-level NMT on a number of datasets. However, document-level NMT is still not widely adopted in real-world translation systems mainly due to the lack of large-scale general-domain training data for document-level NMT. We examine the effectiveness of using Paracrawl  for 
learning document-level translation. Paracrawl is a large-scale parallel corpus crawled from the Internet and contains data from various domains. 
The
 official  Paracrawl corpus was released as    parallel sentences (extracted from parallel webpages) and therefore previous works   only used Paracrawl   for learning sentence-level translation.
 In this work, we  extract parallel paragraphs from Paracrawl parallel webpages  using automatic sentence alignments and  we use the extracted parallel paragraphs as parallel documents for training document-level translation models.
  We show that
document-level NMT models trained with only parallel paragraphs from Paracrawl  can be used to translate real documents   from  TED, News and Europarl, outperforming sentence-level NMT models. We also perform a targeted pronoun evaluation  and show that document-level models trained with Paracrawl data can help  context-aware pronoun
translation.   
We release our
data and code here\footnote{{https://github.com/Yusser96/Exploring-Paracrawl-for-Document-level-Neural-Machine-Translation}}.
\end{abstract}

\section{Introduction}

The Transformer translation model \cite{vaswani2017attention}, which performs sentence-level translation based on attention networks, has achieved great success and significantly improved the state-of-the-art in machine translation. Compared to sentence-level translation, document-level translation \cite{xu-etal-2021-document-graph,bao-etal-2021-g,jauregi-unanue-etal-2020-leveraging,ma-etal-2020-simple,maruf-etal-2019-selective,tu-etal-2018-learning,maruf-haffari-2018-document}
performs translation at document-level and can potentially further improve translation quality, e.g., document-level context can help word disambiguation for translating words with multiple senses, document-level translation can help pronoun translation which requires context outside of the current sentence \cite{muller-etal-2018-large}, document-level translation can improve document-level 
lexical cohesion in the translation \cite{voita-etal-2019-good}.

Document-level neural machine translation (NMT) has received much attention in recent years \cite{bao-etal-2021-g,donato-etal-2021-diverse,fernandes-etal-2021-measuring,kang-etal-2020-dynamic,saunders-etal-2020-using,yu-etal-2020-better,ijcai2020p551,yang-etal-2019-enhancing,kuang-etal-2018-modeling,bawden-etal-2018-evaluating, zhang-etal-2018-improving, voita-etal-2018-context, kuang-xiong-2018-fusing}. Existing works showed that document-level translation can outperform sentence-level translation for a number of datasets, such as TED, News,  Europarl \cite{bao-etal-2021-g,donato-etal-2021-diverse,xu-etal-2021-document-graph}. 
Although document-level NMT has shown promising results on a number of benchmarks, document-level NMT is still not widely adopted in real-world translation systems mainly due to the lack of large-scale general domain training data for document-level NMT.

We examine the effectiveness of using Paracrawl \cite{ banon-etal-2020-paracrawl}  for learning document-level NMT. Paracrawl is a large-scale parallel corpus crawled from the Internet and contains data from various domains. The official   Paracrawl corpus\footnote{{https://paracrawl.eu/}} was released as   parallel sentences (extracted from parallel webpages) and therefore previous works   only used Paracrawl   for learning sentence-level translation. 
In this work, we  extract
parallel paragraphs from  Paracrawl parallel webpages using automatic sentence alignments and we use the extracted parallel paragraphs as parallel documents for
 training document-level translation models.   We show that
document-level NMT models trained with only parallel paragraphs from Paracrawl  can be used to  translate real documents from TED, News and Europarl, outperforming sentence-level NMT models and also improving context-aware pronoun
translation \cite{muller-etal-2018-large}.



\section{Extracting Parallel Paragraphs from Paracrawl}
\label{ddata}
Paracrawl \cite{banon-etal-2020-paracrawl} is a large-scale  parallel corpus crawled from the Internet and   contains texts from various domains. The official   Paracrawl corpus was released as   aligned sentence pairs. For high-resource language pairs like German-English, the Paracrawl corpus provides 278M aligned sentences for the translation task. As Paracrawl was released as aligned sentence pairs, previous work only used Paracrawl for learning sentence-level translation. However, as described in the Paracrawl documentation, Paracrawl was constructed by first identifying aligned webpages (URLs) and then aligning sentences within aligned webpages. Although texts in parallel webpages tend to be noisy and
are   only loosely aligned,
we demonstrate that extracting parallel paragraphs  from Paracrawl parallel webpages using automatic sentence alignments can provide effective training data for learning  document-level translation.
Below we describe how we extract parallel paragraphs from Paracrawl.

\paragraph{Extracting} We extract parallel paragraphs from Paracrawl and examine      the effectiveness of using these aligned paragraphs    for learning document-level NMT.
We used German-English, a high-resource language pair, in our experiments.  As the main purpose of the original Paracrawl project is to collect aligned sentences, Paracrawl did not officially release all the webpage-level aligned texts. Paracrawl only released a subset of all the aligned webpage texts,   for the German-English language pair\footnote{{https://www.statmt.org/paracrawl-benchmarks/}}.  We   extract parallel paragraphs from
the released parallel  webpages:
we first download aligned webpages\footnote{https://www.statmt.org/paracrawl-benchmarks/paracrawl-benchmark.en-de.aligned-docs.xz} (one German webpage is aligned with one English webpage) and then download the automatic sentence alignments (vecalign\footnote{https://www.statmt.org/paracrawl-benchmarks/paracrawl-benchmark.en-de.vecalign.xz})  provided by Paracrawl; we then split a webpage into   paragraphs according to the newline symbol; then  extract aligned paragraphs from aligned webpages according to the automatic sentence alignments. 
We consider two paragraphs to be aligned if sentences in these two paragraphs are aligned to each other and not aligned to any other paragraphs.  We discarded   sentences that are not   one-to-one aligned (e.g. one English sentence aligned to two German sentences) and discarded  sentences that are not aligned to any other sentences. We discarded repeated paragraphs and   discarded paragraphs that only contain a single sentence. Paragraphs with non-monotonic sentence alignments were also discarded.

\begin{table}\small
    \centering
    \begin{tabular}{c|c|c|c|c}
    \hline
    &&&\multicolumn{2}{c}{Words}\\
    &Paragraphs  & Sentences  &En&De \\
    \hline
        Train & 1.5M&5.5M &118M&109M\\
         Dev&402&1504&32K&29K\\
         Test&411&1510&33K&30K\\
         \hline
    \end{tabular}
    \caption{Statistics of parallel paragraphs  extracted from Paracrawl. }
    \label{paradata}
\end{table}

 \begin{table}[t]\small
     \centering
     \begin{tabular}{l|l}
     \hline

     \hline
        Original vecalign   & 147M \\
          After parallel paragraph extraction   & 11.7M \\
          After cleaning & 5.5M\\
        \hline
     \end{tabular}
     \caption{Numbers of remaining sentence pairs after different processing steps.}
     \label{remaindata}
 \end{table}

\begin{table}[t]\small
    \centering
    \begin{tabular}{c|c}
    \hline
     Sentences &  Distribution\\

    \hline
        2 & 34.63\% \\
         3&29.31\%\\
         4&15.99\%\\
         
         5$\sim$10&18.29\%\\
         $>$10&1.77\%\\
         \hline
    \end{tabular}
    \caption{Distribution of paragraph length. For example, the first line of numbers mean 34.63\% of the extracted paragraphs contain 2 sentences.}
    \label{slpara}
\end{table}

\paragraph{Cleaning}
To improve the quality of the extracted parallel paragraphs, we    removed sentences which do not belong to the correct language\footnote{https://pypi.org/project/langid/},  removed paragraphs that are too short (contain less than 30 words) and removed paragraphs with more than 50\% overlap. 

 We
randomly split the extracted parallel paragraphs into training, development and test sets as shown in Table~\ref{paradata}. 
Note that, compared to the officially released Paracrawl corpus, the size of parallel paragraphs that we extracted
from Paracrawl is still relatively small. This is because (i) the aligned webpages provided by Paracrawl we used for parallel paragraph extraction is only a subset of all Paracrawl data (ii) many parallel sentences were discarded due to our strict extraction rules as shown in Table~\ref{remaindata}. For future work, we will collect more webpage-level aligned Paracrawl data and test more flexible extraction rules.

 We show the length  statistics  of the parallel paragraphs that we  extracted from Paracrawl in Table~\ref{slpara}. 
  Compared to normal documents, the parallel paragraphs extracted from Paracrawl are generally much shorter. However, sentences in the same paragraph are usually closely related and can  provide useful context information to help the translation of each other. In our experiments, we use   the extracted parallel   paragraphs as parallel documents to train document-level NMT models and we show that  document-level NMT models trained with only parallel paragraphs from Paracrawl  can be used to translate real documents  from TED, News and Europarl, outperforming sentence-level NMT models.

 \section{Document-level Translation with Paracrawl}
\label{expe}
\subsection{Modeling}
Previous works have shown that   document-level NMT models can outperform sentence-level models on several benchmarks.
We adopted the recent document-level translation model, 
G-Transformer  \cite{bao-etal-2021-g}\footnote{{https://github.com/baoguangsheng/g-transformer}}, in our experiments to test the effectiveness of using Paracrawl data for learning document-level translation.
The G-Transformer model is based on the standard sentence-level Transformer \cite{vaswani2017attention}, but uses a whole document together as the input of the model and then generates translation for the whole document. G-Transformer improves the standard Transformer model for document-level translation with extra group tags and group attention. Each word (both source and target words) in the document is assigned with a group tag to indicate which sentence this word belongs to. Compared to the standard Transformer attention, G-Transformer computes group attention using group tags   
       to encourage local attention and  reducing the hypothesis space of the
attention, especially from target to source, for long documents. 
 G-Transformer outperformed sentence-level Transformer and obtained  new
state-of-the-art results
   on three document-level translation
benchmarks.

\subsection{Experimental Setting}

 We tokenize and truecase all data
with MOSES \citep{koehn-etal-2007-moses} scripts, and then perform subword segmentation with
  byte pair encoding   (BPE) \citep{sennrich-etal-2016-neural} using 30k merging
operations.
For both the sentence-level Transformer and the document-level G-Transformer, we used the base model setting \cite{vaswani2017attention} with 6-layer encoder/decoder, 512-dimension word embedding and 2048 hidden units for the feed forward networks.  
Following the G-Transformer experimental settings \cite{bao-etal-2021-g},
 we set the max length of a document to     512 BPE tokens (if a document is longer than 512 tokens, we split it into multiple 
  instances).
  For model training, we first pretrained a standard sentence-level Transformer for 100k training steps and then finetuned the sentence-level Transformer to learn document-level G-Transformer for another 100k training steps. For a fair comparison between sentence-level and document-level translation, we also applied pretraining for sentence-level translation, i.e. we train a sentence-level Transformer with 100k pretraining and 100k finetuning 
 steps in order to compare with G-Transformer.

 \begin{table*}[]\small
     \centering
     \begin{tabular}{ll|ll|ll|ll|ll}
     \hline
    & &\multicolumn{4}{|c|}{Train-Paracrawl-Only}&\multicolumn{4}{|c}{Train-Combined  }\\
     
     &&EnDe&&DeEn&&EnDe&&DeEn\\
      &    &sent &doc&sent&doc&sent&doc&sent&doc \\
      \hline
       BLEU&   Paracrawl& 26.57&\textbf{26.96} &30.94&\textbf{31.94}$^\dagger$ &26.44&\textbf{27.40}$^\dagger$  &31.36&\textbf{31.92}$^\dagger$ \\
        &  Europarl&22.78&\textbf{23.04} &27.76&\textbf{28.69}$^\dagger$ &28.60&\textbf{29.30}$^\dagger$ &35.53&\textbf{35.90}$^\dagger$ \\
         & TED&23.56&\textbf{24.39}$^\dagger$ &27.89&\textbf{28.24} &25.70&\textbf{26.13} &30.15&\textbf{31.49}$^\dagger$ \\
          &News&31.55&\textbf{32.07}$^\dagger$ &34.59&\textbf{35.39}$^\dagger$ &33.08&\textbf{33.89}$^\dagger$ &35.99&\textbf{36.39}\\
          \hline
          COMET&Paracrawl&16.16&\textbf{17.99} &18.90&\textbf{20.99}$^\dagger$ &17.66&\textbf{20.44}$^\dagger$ &20.60&\textbf{22.02}\\
          &Europarl&38.34&\textbf{39.84}$^\dagger$ &42.39&\textbf{45.10}$^\dagger$ &54.10&\textbf{54.80}$^\dagger$ &55.78&\textbf{56.31}$^\dagger$\\
          &TED&23.14&\textbf{23.77} &41.34&\textbf{42.62}&34.21&\textbf{35.25}&46.02&\textbf{46.92}\\
          &News&33.84&\textbf{35.51}$^\dagger$ &44.45&\textbf{48.26}$^\dagger$ &42.49&\textbf{44.20}$^\dagger$ &49.66&\textbf{51.35}$^\dagger$\\
          \hline
     \end{tabular}
     \caption{Translation results.  Train-Paracrawl-Only:  only Paracrawl as training data. Train-Combined: Paracrawl, Europarl, TED and News combined as training data. $^\dagger$ represents a significant difference at the p <
0.01  level.}
     \label{mainres}
 \end{table*}

\begin{table*}[t!]\small
\centering
\begin{tabular}{l|llll|llll}
\hline
&\multicolumn{4}{c}{Train-Paracrawl-only}&\multicolumn{4}{|c}{Train-Combined}\\
 &total &es &er &sie &total &es &er &sie \\
\hline
sent &0.43& 0.91 & 0.15 & 0.24  & 0.43& \textbf{0.94} & 0.17 &0.20 \\ 
doc &\textbf{0.55}& \textbf{0.93}& \textbf{0.34} &\textbf{0.37}  &\textbf{0.60}& 0.92 &\textbf{0.44} &\textbf{0.44}   \\\hline
\end{tabular}

\caption{ Accuracy on contrastive pronoun test set  with regard to reference pronoun. }
\label{parapron}
\end{table*}

\begin{table*}[t!]\small
\centering
\begin{tabular}{l|l|l|l|l}
\hline
&\multicolumn{2}{c|}{Train-Paracrawl-Only}&\multicolumn{2}{c}{Train-Combined}\\
&inside&outside&inside&outside\\
 & current sentence & current sentence &current sentence& current sentence \\
\hline
sent &\textbf{0.68}& 0.37   &0.71 &0.37  \\ 
doc &0.67&\textbf{0.52}  &\textbf{0.76} &\textbf{0.56}   \\\hline
\end{tabular}

\caption{ Accuracy on contrastive pronoun test set with regard to antecedent location. }
\label{paraloca}
\end{table*}

 \subsection{Evaluating with BLEU, COMET  and Targeted Pronoun Evaluation} We used BLEU \cite{papineni2002bleu} and COMET \cite{rei-etal-2020-comet} for translation quality evaluation\footnote{Both BLEU and COMET scores were computed on sentence-level using\\ https://github.com/mjpost/sacrebleu   \\ https://unbabel.github.io/COMET/html/index.html}
and  we performed significance testing   with bootstrap resampling \cite{koehn-2004-statistical}.
In addition to BLEU and COMET which evaluate the general translation quality, we also performed a targeted evaluation for pronoun translation following \citet{muller-etal-2018-large}'s work. The correct translation of a pronoun often requires context outside   of the current sentence. Therefore, evaluation of pronoun translation can demonstrate the advantage of document-level NMT models. Following \citet{muller-etal-2018-large}'s work, we 
computed the accuracy of our models choosing the correct translation for the English pronoun ``it" from the three possible German words ``es", ``er" and ``sie". We used a context of 5 sentences for the test data\footnote{https://github.com/ZurichNLP/ContraPro} provided by \citet{muller-etal-2018-large} for pronoun evaluation.

\begin{table}[]\small
    \centering
    \begin{tabular}{l|l|l|l|l}
    \hline
         & \multicolumn{2}{c|}{T-P-O}&\multicolumn{2}{c}{T-C} \\
         &EnDe&DeEn&EnDe&DeEn\\
         \hline
         Paracrawl&26.88&31.37&27.35&31.65\\
         Europarl&23.22&28.47&29.03&35.61\\
         TED&24.24&28.04&25.62&29.91\\
         News&31.86&34.92&33.41&35.99\\
         \hline
    \end{tabular}
    \caption{Translation results (BLEU) of using the document-level G-Transformer for translating single sentences. T-P-O: Train-Paracrawl-Only. T-C: Train-Combined.}
    \label{bleusingle}
\end{table}

\subsection{Training with Only Paracrawl}
\label{seconlypara}
We trained both sentence-level Transformers and document-level G-Transformers
 with only Paracrawl  (see Table~\ref{paradata}) as training data. Parallel paragraphs were used as parallel documents to train document-level G-Transformers and parallel sentences contained in parallel paragraphs were used to train sentence-level Transformers.\footnote{Therefore,  training data for  the document-level model and the sentence-level model contain the same number of sentence pairs.}
For evaluation, we used test data from
 4 datasets, Paracrawl, Europarl, TED and News. For Europarl, TED and News, we used the same test data following the original G-Transformer  \cite{bao-etal-2021-g} work, i.e., the Europarl, TED and News test sets are parallel documents in contrast to the Paracrawl test set which is parallel paragraphs.
 BLEU and COMET scores are given in Table~\ref{mainres} as Train-Paracrawl-Only.
\begin{table}[t!]\small
    \centering
    \begin{tabular}{l|l|l|l}
    \hline
         \multicolumn{2}{c|}{T-P-O}&\multicolumn{2}{c}{T-C}  \\
         inside&outside&inside&outside \\
         \hline
         0.67&0.36&0.78&0.35\\
         \hline
    \end{tabular}
    \caption{Accuracy on contrastive pronoun test set with regard to antecedent location when using the document-level G-Transformer for translating single sentences without context.}
    \label{prounsingle}
\end{table}
  The document-level G-Transformers  achieved higher translation quality than sentence-level Transformers for all 4 test sets and only used Paracrawl as training data, which demonstrates that Paracrawl can provide useful document-level information  for effective training of document-level NMT models. 
Table~\ref{mainres} also shows that the document-level information contained in Paracrawl is robust across domains as    
 document-level G-Transformers trained with only  Paracrawl data can help to    translate real documents from  TED, News and Europarl test sets.

Table~\ref{parapron} and Table~\ref{paraloca} give results of targeted pronoun evaluation. Results show that the document-level model, trained with only Paracrawl data, significantly outperformed the sentence-level model for pronoun translation especially when the antecedent location of a pronoun is outside of the current sentence (see Table~\ref{paraloca}), which again demonstrates  that Paracrawl can provide useful information outside of the current sentence for effective learning of document-level NMT.

\subsection{Training with Paracrawl, TED, News and Europarl Combined}

We also trained translation models with training data from Paracrawl, Europarl, TED and News combined. The sentence-level Transformers were trained with sentence pairs from all the 4 training datasets. The document-level G-Transformers were trained with parallel paragraphs from Paracrawl and parallel documents from Europarl, TED and News. 
 We then computed BLEU     and COMET scores for the 4 test sets as shown in Table~\ref{mainres} as Train-Combined. The targeted pronoun evaluation results  are also given in Table~\ref{parapron} and Table~\ref{paraloca} as {Train-Combined}.
 Results show that using additional parallel documents together with parallel (Paracrawl) paragraphs as training data   further improved the general translation quality (BLEU and COMET) and also helped the document-level model to obtain a higher accuracy for targeted pronoun translation evaluation.

\subsection{Document-level G-Transformer for Translating Single Sentences}
We also evaluated how the document-level G-Transformer (trained with parallel documents) performs for translating single sentences (i.e. each sentence is considered as a single document  at test time). Table~\ref{bleusingle} shows that, when considering the general translation quality (e.g. BLEU),  document-level G-Transformers can perform well for translating single sentences, even outperforming the sentence-level Transformers in Table~\ref{mainres} for most of the test sets. 
However, the targeted pronoun evaluation results in Table~\ref{prounsingle} show that the accuracy of translating pronouns with antecedent location outside of the current sentence dropped more than 10\% compared to the document-level model in Table~\ref{paraloca}, which demonstrates that document-level G-Transformers indeed  require context outside of the current sentence for improving pronoun translation.


\section{Conclusion      }
 \label{concl}
As document-level translation lacks large-scale general-domain document-level training data, we   examine the effectiveness of using Paracrawl data for learning document-level translation.
 Paracrawl was officially released as parallel sentences extracted from parallel webpages.
In this work, we  extract parallel paragraphs from  Paracrawl aligned webpages   using  automatic sentence alignments 
and we use the extracted parallel paragraphs as parallel documents for training document-level NMT models. We show that document-level  
models trained with only parallel paragraphs from Paracrawl  can be used to translate real documents from
 TED, News and Europarl, outperforming sentence-level Transformers and also improving context-aware pronoun
translation.

 \section*{Limitations}

 Compared to the officially released Paracrawl corpus (278M sentence pairs for German-English), the size of parallel paragraphs  that we extracted from Paracrawl is still relatively small. This is because (i) Paracrawl only released a subset of all webpage-level aligned  texts and we only extracted parallel paragraphs from these released webpage texts (ii) we used very strict rules for extracting parallel paragraphs from Paracrawl and many parallel sentences were discarded by our extraction rules. For future work, we will  collect more Paracrawl webpage-level aligned data for parallel paragraph extraction and we will test more flexible extraction rules.

\section*{Acknowledgements}

The authors acknowledge the financial support by the German Federal Ministry for Education and Research (BMBF) within the projects ``CORA4NLP" {01IW20010} and  ``KI-Servicezentrum Berlin Brandenburg" {01IS22092}.

\bibliography{anthology,custom}
\bibliographystyle{acl_natbib}

\appendix

\end{document}